for a reaction of nth order the ratio of t(half)/t(3/4) is a function of 


 

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
Therate laworrate equationfor achemical reactionis an equation that links thereaction ratewith concentrations or pressures of reactants and constant parameters To determine the rate equation for a particular system one combines the reaction rate with amass balancefor the system.

The half-life for a general nth order reaction is given by the expression .
t(1/2) = {2^(n-1) - 1} / {k(n-1)([A]^(n-1))}
the integrated rate law will be given by
1/A^n-1=1/A0^n-1 +(n-1)kt
therefore t1/2 /t3/4=2^n-1 -1/4^n-1 -1
Prabhleen Kaur
6 years ago
t1/2=1/kn*1/(n-1)*1/(a^n-1)*(2^(n-1)-1)
t3/4=1/kn*1/(n-1)*1/(a^n-1)*(4^(n-1)-1)
Ratio 
t1/2/t3/4=(2^n-1-1)/(4^n-1-1)